# HoSNNs: Adversarially-Robust Homeostatic Spiking Neural Networks with Adaptive Firing Thresholds

### v3

@@ A.1 LIF Dynamics
A.2 Neural Dynamic Signature
@@
To derive the dynamics of NDS, we start by revisiting Equation (15), rewriting it with respect to $x$
@@
We take this approximation to address discontinuity of Dirac funtion, as $r_{i}(x)$ and $s_{i}(t|x)$ have the same integral value over time: the number of neuron firings. Substituting Equation (21) into Equation (20) and computing the expectation on both sides, we have:
@@
With the definition from Equation (19), the dynamics of NDS can be expressed as:
@@ A.3 TA-LIF Dynamics
The synaptic input $I_{i}(t)$, the spike generation function $s_{i}(t)$ and post-synaptic current dynamics of TA-LIF are defined same as (16)(17)(18). For a specific network parameter $\bm{W}$ and a sample $x^{\prime}$ drawn from $\mathcal{D^{\prime}}$, the dynamic equation governing the threshold $V^{i}_{th}(t)$ is:
@@
where the error signal, utilizing the NDS as given in (19), is defined as:
@@
Applying the continuity approximation for the Dirac function as per (21) and incorporating the conditional dependency of $\bm{W}$ and $x^{\prime}$, rewriting the dynamics for TA-LIF (27) as:
@@
Subtracting (26) from (30) and employing (29), denoting
@@
Differentiating (32) with respect to time and utilizing the threshold dynamics from (29), we obtain:
@@
For succinctness, we will omit dependencies on $\bm{W}$ and $x^{\prime}$, resulting in TA-LIF dynamics in the main text (25):
@@
Appendix B Dynamic Stability Analysis
In this section, we analyze the stability of (34) and (35) to explore the influence of our dynamic threshold mechanism on the noise suppression ability of TA-LIF neuron.
B.1 BIBO Stability of Equation (34)
Characteristic Equation: We first show the BIBO (Bounded Input, Bounded Output) stability Ogata [2010] of TA-LIF neurons based on (34). The characteristic equation of (34) of non-silent ($r_{i}>0$) and non-degenerating ($\tau_{m},\theta_{i}>0$) TA-LIF neurons are:
@@ B.1 BIBO Stability of Equation (34)
For a second-order system to be BIBO, the roots of its characteristic equation must be a negative real or have a negative parts, which is clearly the case for the TA-LIF model under the above three situations, affirming the BIBO stability of (34). The BIBO stability signifies that with the bounded driving input to system (34), the deviation of the TA-LIF neuron’s membrane potential from its targeted NDS is also bounded, demonstrating the well control of the growth of error $e_{i}(t)$.
B.2 Stability of Equation (34) Under White Noise
To elucidate the dynamic characteristics of TA-LIF further, we adopt the prevalent method Abbott and Van Vreeswijk [1993], Brunel [2000], Gerstner et al. [2014], Renart et al. [2004], approximating $\Delta I(t)$ with a Wiener process. This approximation effectively represents small, independent, and random perturbations. Hence, the driving force in equation (34) $\frac{d\Delta I_{i}(t)}{dt}$ can be modeled by a Gaussian white noise $F(t)$, leading to the well-established Langevin equation in stochastic differential equations theory Kloeden et al. [1992], Van Kampen [1992], Risken [1996]:
@@ B.2 Stability of Equation (34) Under White Noise
where the sum has to be taken over all the different ways in which one can divide the $2n$ time points $t_{1}\cdots t_{2n}$ into $n$ pairs. Under this assumption (39), the solution of the Langevin equation(38) is Uhlenbeck and Ornstein [1930], Wang and Uhlenbeck [1945]:
@@
where $\Delta e_{i}(t)=e_{i}(t)-\langle e_{i}(t)\rangle$ and $\omega_{1}=\sqrt{{r_{i}}{\theta_{i}}-\frac{1}{4\tau^{2}_{m}}}$. While under the same assumptions (39), equation (35) yields:
@@
Obviously, gaussian white noise with zero mean (39) leads $\langle e_{i}(t)\rangle=0$, $\langle[\Delta e_{i}(t)]^{2}\rangle=\langle e_{i}^{2}(t)\rangle$. Hence,
@@
Appendix C Experiment Setting Details
@@
Here, the notation 32C3 represents a convolutional layer with 32 filters of size $3\times 3$, and P stands for a pooling layer using $2\times 2$ filters. For the CIFAR10 and CIFAR100 datasets, we incorporated Batch normalization layers and dropout mechanisms to mitigate overfitting and elevate the performance of the deep networks. In our experiments with FashionMNIST, SVHN and CIFAR10, the output spike train of LIF neurons was retained to compute the kernel loss, as described in Zhang and Li [2020]. For CIFAR100, we directly employed softmax for performance.
@@
For all HoSNN experiments, a preliminary training phase was carried out using an LIF SNN, sharing the same architecture, on the clean datasets to deduce the NDS. Hyperparameters for LIF and TA-LIF neurons included a simulation time $T=5$ , a Membrane Voltage Constant $\tau_{m}=5$, and a Synapse Constant $\tau_{s}=3$. For the TA-LIF results in the main text, we assigned $\theta_{i}$ initialization values of 5 for FashionMNIST, SVHN, CIFAR10 and 3 for CIFAR100. All neurons began with an initial threshold of 1. The step function was approximated using $\sigma(x)=\frac{1}{1+e^{-5x}}$, where $x=u(t)-V_{th}(t)$ and the BPTT learning algorithm was employed. For TA-ALIF neurons, the learning rate for $\theta_{i}$ was set at $1/10$ of the rate designated for weights, ensuring hyperparameter stability during training. We also constrained $\theta_{i}$ to remain non-negative during optimization, ensuring a possible transition from TA-LIF to LIF. For the generation of all gradient-based adversarial attacks, we assume that the attacker can know all $V_{th}(t)$ and use it in the gradient, even if it is generated dynamically and is not stored as network parameters. During training we use $V_{th}(t)=V_{th}(0)$. We utilized the Adam optimizer with hyperparameters betas set to (0.9, 0.999), and the $lr=5\times 10^{-4}$ with cosine annealing learning rate scheduler ($T=$ epochs). We set batch size to 64 and trained for 200 epochs. All images were transformed into currents to serve as network input. Our code is adapted from Zhang and Li [2020]. The experiment used four NVIDIA A100 GPUs. For cifar10 and cifar100, it took up to about 48 hours for adversarial training.
@@
Regarding adversarial attack, we use an array of attack strategies, including FGSM, RFGSM, PGD, and BIM. For both CIFAR10 and CIFAR100, we allocated an attack budget with $\epsilon=8/255$. For iterative schemes like PGD, we set $\alpha=\epsilon/3$ and $steps=7$, aligning with the recommendations in Ding et al. [2022]. For the adversarial training phase, FGSM training was used with $\epsilon$ values of 2/255 for CIFAR10 (as per Ding et al. [2022]) and 4/255 for CIFAR100, following Kundu et al. [2021].
Appendix D Checklist for Gradient obfuscation
Obfuscated gradients are a type of gradient masking that lead to a false sense of security when defending against adversarial examples Athalye et al. [2018], Carlini et al. [2019]. Here we perform sanity checks including three obfuscated types and checklist as per Athalye et al. [2018].
First we examine three types of obfuscated gradients. Specifically, we use the same surrogate to train HoSNN from scratch and deliver attacks, decent clean accuracy and smooth training process indicate that it’s not Shattered Gradient with nonexistent or incorrect value. As our defense doesn’t introduce any random factors, Stochastic Gradient is not applicable. Our method also doesn’t include any multiple iterations of neural network evaluation, so Vanishing/Exploding Gradient are also not applicable.
@@ Appendix D Checklist for Gradient obfuscation
(4) Unbounded attacks can’t reach $\sim 100\%$ success
@@
Then let’s check the item in checklist one by one.
(a) Fashion-MNIST
(b) SVHN
(c) CIFAR-10
(d) CIFAR-100
Figure 7: For Test (1). Performance of HoSNN under white-box FGSM and PGD attack.
For Test (1): Single-step attack performs better compared to iterative attacks, we plot the curves of white-box FGSM and PGD attacks on four datasets in Figure 7, with attack budgets $\epsilon$ from 0 to 64/255 to ensure that the network can be completely fooled. The red curve is the accuracy under FGSM, while the blue curve is PGD. From the Figure 7 we can confirm that all iterative attacks are much stronger than single-step attack.
(a) Fashion-MNIST
(b) SVHN
(c) CIFAR-10
(d) CIFAR-100
Figure 8: For Test (2). Performance of HoSNN under white-box and black-box and PGD attack.
For Test (2): Black-box attacks performs better compared to white-box attacks, we plot the curves of white-box and black-box PGD attacks on four datasets in Figure 8, with attack budgets $\epsilon$ from 0 to 64/255 to ensure that the network can be completely fooled. The green curve is the accuracy under black-box PGD attack, while the blue curve is under white-box. From the Figure 8 we can confirm that all white-box attacks are much stronger than black-box.
(a) Fashion-MNIST
(b) SVHN
(c) CIFAR-10
(d) CIFAR-100
Figure 9: For Test(3) and (4). Performance of HoSNN under larger white-box PGD attack.
For Test (3) & (4): Increasing perturbation bound can’t increase attack strength & Unbounded attacks can’t reach $\sim 100\%$ success, we plot the curves of white-box and black-box PGD attacks on four datasets in Figure 9, with attack budgets $\epsilon$ from 0 to 64/255 to ensure that the network can be completely fooled. The blue curve is HoSNN’s accuracy under white-box PGD attack, while the yellow curve is SNN’s baseline. From the Figure 9 we can confirm that as perturbation bound increasing HoSNN’s accuracy is decreasing and all unbounded attacks reach $\sim 100\%$ success.
For Test (5) Adversarial example can be found through random sampling, since all gradient-based attacks work, there is no need to use random sampling methods. Therefore Test (5) passed obviously. All five gradient obfuscation test passed.
Appendix E Supplementary Data
@@
Table 4: HoSNN accuracy data for Test3. We showed the performance of SNN and HoSNN under white-box PGD attack. Our data shows that increasing perturbation bound can increase attack strength and the accuracy can drop to 0 as the attack becomes stronger.

### v4

@@ B.1 LIF Dynamics
B.2 Neural Dynamic Signature
@@
To derive the dynamics of NDS, we start by revisiting Equation equation 16, rewriting it with respect to $x$
@@
We take this approximation to address discontinuity of Dirac funtion, as $r_{i}(x)$ and $s_{i}(t|x)$ have the same integral value over time: the number of neuron firings. Substituting Equation equation 22 into Equation equation 21 and computing the expectation on both sides, we have:
@@
With the definition from Equation equation 20, the dynamics of NDS can be expressed as:
@@ B.3 TA-LIF Dynamics
The synaptic input $I_{i}(t)$, the spike generation function $s_{i}(t)$ and post-synaptic current dynamics of TA-LIF are defined same as equation 17equation 18equation 19. For a specific network parameter $\bm{W}$ and a sample $x^{\prime}$ drawn from $\mathcal{D^{\prime}}$, the dynamic equation governing the threshold $V^{i}_{th}(t)$ is:
@@
where the error signal, utilizing the NDS as given in equation 20, is defined as:
@@
Applying the continuity approximation for the Dirac function as per equation 22 and incorporating the conditional dependency of $\bm{W}$ and $x^{\prime}$, rewriting the dynamics for TA-LIF equation 28 as:
@@
Subtracting equation 27 from equation 31 and employing equation 30, denoting
@@
Differentiating equation 33 with respect to time and utilizing the threshold dynamics from equation 30, we obtain:
@@
For succinctness, we will omit dependencies on $\bm{W}$ and $x^{\prime}$, resulting in TA-LIF dynamics in the main text equation 26:
@@
Appendix C Dynamic Stability Analysis
In this section, we analyze the stability of equation 35 and equation 36 to explore the influence of our dynamic threshold mechanism on the noise suppression ability of the TA-LIF neuron.
C.1 BIBO Stability of Equation equation 35
Characteristic Equation: We first show the BIBO (Bounded Input, Bounded Output) stability Ogata (2010) of TA-LIF neurons based on equation 35. The characteristic equation of equation 35 of non-silent ($r_{i}>0$) and non-degenerating ($\tau_{m},\theta_{i}>0$) TA-LIF neurons is:
@@ C.1 BIBO Stability of Equation equation 35
For a second-order system to be BIBO, the roots of its characteristic equation must be negative real or have negative parts, which is clearly the case for the TA-LIF model under the above three situations, affirming the BIBO stability of equation 35. The BIBO stability signifies that with the bounded driving input to system equation 35, the deviation of the TA-LIF neuron’s membrane potential from its targeted NDS is also bounded, demonstrating the well control of the growth of error $e_{i}(t)$.
C.2 Stability of Equation equation 35 Under White Noise
To elucidate the dynamic characteristics of TA-LIF further, we adopt the prevalent method Abbott & Van Vreeswijk (1993); Brunel (2000); Gerstner et al. (2014); Renart et al. (2004), approximating $\Delta I(t)$ with a Wiener process. This approximation effectively represents small, independent, and random perturbations. Hence, the driving force in equation equation 35 $\frac{d\Delta I_{i}(t)}{dt}$ can be modeled by a Gaussian white noise $F(t)$, leading to the well-established Langevin equation in stochastic differential equations theory Kloeden et al. (1992); Van Kampen (1992); Risken (1996):
@@ C.2 Stability of Equation equation 35 Under White Noise
where the sum has to be taken over all the different ways in which one can divide the $2n$ time points $t_{1}\cdots t_{2n}$ into $n$ pairs. Under this assumption equation 40, the solution of the Langevin equationequation 39 is Uhlenbeck & Ornstein (1930); Wang & Uhlenbeck (1945):
@@
where $\Delta e_{i}(t)=e_{i}(t)-\langle e_{i}(t)\rangle$ and $\omega_{1}=\sqrt{{r_{i}}{\theta_{i}}-\frac{1}{4\tau^{2}_{m}}}$. While under the same assumptions equation 40, equation equation 36 yields:
@@
Obviously, Gaussian white noise with zero mean equation 40 leads $\langle e_{i}(t)\rangle=0$, $\langle[\Delta e_{i}(t)]^{2}\rangle=\langle e_{i}^{2}(t)\rangle$. Hence,
@@
C.3 A more intuitive explanation
We would like to provide a more intuitive explanation of Eq 9 and Eq 10. Notice that when $\theta_{i}=0$ according to Eq 10, TA-LIF no longer has a dynamically changing threshold and thus degenerates into standard LIF. Its second-order dynamics degenerates into:
$\tau_{m}\frac{d^{2}e_{i}(t)}{dt^{2}}+\frac{de_{i}(t)}{dt}=\varepsilon(t)$
An example to intuitively understand these two formulas is that they correspond to a driven damped oscillator in the physical world, that is, the movement of a ball connected to a spring under a driving force and damping environment. Considering $e_{i}(t)$ as the distance the ball deviates from the equilibrium position, we can understand the physical meaning of each term in the formula: $\tau_{m}\frac{d^{2}e_{i}(t)}{dt^{2}}$ describes the acceleration of the ball; $\frac{de_{i}(t)}{dt}$ describes the frictional resistance of the ball; $r_{i}\tau_{m}\theta_{i}e_{i}(t)$ describes the restoring force of the spring, which is the tendency of the spring to pull the ball toward the equilibrium position ($e_{i}(t)=0$); $\varepsilon(t)$ describes the driving force exerted by the environment on the ball.
With this explanation, we can easily understand the difference between the second-order dynamics of TA-LIF and LIF. The key term is $r_{i}\tau_{m}\theta_{i}e_{i}(t)$. For TA-LIF, it is like a small ball that is constantly pulled toward its initial equilibrium position by a spring when it is subjected to external disturbances: the result is that the ball will not deviate too far ($e_{i}(t)$ remains small). However, there is no such term in the dynamics of LIF, which causes $e_{i}(t)$ to become larger under external disturbances.
Appendix D Experiment Setting Details
@@
Here, the notation 32C3 represents a convolutional layer with 32 filters of size $3\times 3$, and P stands for a pooling layer using $2\times 2$ filters. For the CIFAR10 and CIFAR100 datasets, we incorporated batch normalization layers and dropout mechanisms to mitigate overfitting and elevate the performance of the deep networks. In our experiments with FashionMNIST, SVHN, and CIFAR10, the output spike train of LIF neurons was retained to compute the kernel loss, as described in Zhang & Li (2020). For CIFAR100, we directly employed softmax for performance.
@@
For all HoSNN experiments, a preliminary training phase was carried out using an LIF SNN, sharing the same architecture, on the clean datasets to deduce the NDS. Hyperparameters for LIF and TA-LIF neurons included a simulation time $T=5$ , a Membrane Voltage Constant $\tau_{m}=5$, and a Synapse Constant $\tau_{s}=3$. For the TA-LIF results in the main text, we assigned $\theta_{i}$ initialization values of 5 for FashionMNIST, SVHN, CIFAR10 and 3 for CIFAR100. All neurons began with an initial threshold of 1. The step function was approximated using $\sigma(x)=\frac{1}{1+e^{-5x}}$, where $x=u(t)-V_{th}(t)$ and the BPTT learning algorithm was employed. For TA-ALIF neurons, the learning rate for $\theta_{i}$ was set at $1/10$ of the rate designated for weights, ensuring hyperparameter stability during training. We also constrained $\theta_{i}$ to remain non-negative during optimization, ensuring a possible transition from TA-LIF to LIF. For the generation of all gradient-based adversarial attacks, we assume that the attacker can know all $V_{th}(t)$ and use it in the gradient, even if it is generated dynamically and is not stored as network parameters. During training, we use $V_{th}(t)=V_{th}(0)$. We utilized the Adam optimizer with hyperparameters betas set to (0.9, 0.999), and the $lr=5\times 10^{-4}$ with cosine annealing learning rate scheduler ($T=$ epochs). We set batch size to 64 and trained for 200 epochs. All images were transformed into currents to serve as network input. Our code is adapted from Zhang & Li (2020). The experiment used four NVIDIA A100 GPUs. For CIFAR10 and CIFAR100, it took up to about 48 hours for adversarial training.
@@
Regarding adversarial attack, we use an array of attack strategies, including FGSM, RFGSM, PGD, and BIM. For both CIFAR10 and CIFAR100, we allocated an attack budget with $\epsilon=8/255$. For iterative schemes like PGD, we set $\alpha=2.5*\epsilon/steps$ and $steps=7,20,40$, aligning with the recommendations in Ding et al. (2022). For the adversarial training phase, FGSM training was used with $\epsilon$ values of 2/255 for CIFAR10 as per Ding et al. (2022) and 4/255 for CIFAR100, following Kundu et al. (2021).
@@
(4) Unbounded attacks can’t reach $100\%$ success
@@
Table 6: Checklist for gradient obfuscation
Appendix E More Experiment Data
Obfuscated gradients are a type of gradient masking that leads to a false sense of security when defending against adversarial examples Athalye et al. (2018); Carlini et al. (2019). Here we perform sanity checks including three obfuscated types and a checklist as per Athalye et al. (2018). First, we examine three types of obfuscated gradients. Specifically, we use the same surrogate to train HoSNN from scratch and deliver attacks; decent clean accuracy and a smooth training process indicate that it’s not Shattered Gradient with nonexistent or incorrect value. As our defense doesn’t introduce any random factors, Stochastic Gradient is not applicable. Our method also doesn’t include any multiple iterations of neural network evaluation, so Vanishing/Exploding Gradient are also not applicable.
Items to identify gradient obfuscation HoSNN Experiment
(1) Single-step attack performs better compared to iterative attacks Fig 5 and Table 8
(2) Black-box attacks perform better compared to white-box attacks Fig 6 and Table 9
(3) Increasing perturbation bound can’t increase attack strength Fig 7 and Table 10
(4) Unbounded attacks can’t reach $100\%$ success Fig 7 and Table 10
(5) Adversarial example can be found through random sampling Fig 8
Table 7: Checklist for gradient obfuscation
@@ Appendix E More Experiment Data
Table 10: HoSNN accuracy data for Test3&4. We showed the performance of SNN and HoSNN under white-box PGD attack. Our data shows that increasing perturbation bound can increase attack strength and the accuracy can drop to 0 as the attack becomes stronger.